OpenStudy (anonymous):

A varies inversely with b. when a is .25 b is 4 fine the value of B when a is 5

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

you have this situation.. \[a = \frac{ 1 }{ b }\] now if a=0.25 then b=4 it is 0.25 =1/4 now if a=5 then.. 5=1/b =>b=1/5=0.2